# Oblique boundary value problems for degenerate Hessian quotient type equations

### v1

@@ 4.1. Global second order estimates
In this subsection, we derive global second order estimates for admissible solutions of problem (1.1).
@@
Let $u\in C^{4}(\Omega)\cap C^{3}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{3}$ domain $\Omega\subset\mathbb{R}^{n}$. Suppose $f^{\frac{1}{k-l}}(x,z)\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$, then there exists a positive constant $C$ depending on $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $\overline{\Omega}$, $n$, $k$, $l$, $\gamma$ such that
@@
Consider an auxiliary function for any $\xi=(\xi^{1},\xi^{2},\cdots,\xi^{n})\in\mathbb{R}^{n}$
@@
Assume that $V$ attains its the interior maximum at $\left(x_{0},\xi_{0}\right)$. We may also suppose $\xi_{0}=(1,0,\cdots,0)$ and the matrix $u_{ij}(x_{0})$ is diagonal with $u_{11}(x_{0})\geq u_{22}(x_{0})\geq\cdots\geq u_{nn}(x_{0})$ under an orthogonal argument. Set
@@
In the rest of the proof, all the calculations are made at $x_{0}$. Since $f^{\frac{1}{k-l}}\in C^{1,1}$ and Lemma 2.5, we obtain
(4.5) $\partial_{11}f\geq-C\sum_{i=1}^{n}F^{ii}+\left(1-\frac{1}{k-l}\right)\frac{(\partial_{1}f)^{2}}{f}.$
By (2.4)-(2.6), (4.4)-(4.5) and Lemma 2.4
(4.6) $\begin{split}0\geq L\widetilde{V}=&\;F^{ij}u_{11ij}+F^{ij}u_{1i}u_{1j}+F^{ij}u_{1}u_{1ij}\\ \geq&\;-C\sum_{i=1}^{n}F^{ii}+u_{11}^{2}\sum_{i=1}^{n}F^{ii},\end{split}$
@@ 4.2. Boundary second order estimates
In this part, we establish boundary second order estimates, and divide the proof into three cases: double tangential derivatives, mixed tangential-oblique derivatives and double oblique derivatives. We assume that the function $G(x,z,p)$ and the unit inner normal vector field $\nu$ have been extended to $\overline{\Omega}$, to be constant along normals to $\partial\Omega$ in some neighbourhood $\Omega_{0}$ of $\partial\Omega$.
@@
Let $u\in C^{4}(\Omega)\cap C^{3}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{3}$ domain $\Omega\subset\mathbb{R}^{n}$. Then for any tangential vector field $\tau$, we have
@@
where $\beta=G_{p}(x,z,p)$ and the positive constant $C$ depends on $G,\overline{\Omega},|u|_{C^{1}}$.
@@
Differentiating $G(x,z,p)=0$ in problem (1.1) with respect to $\tau$, we have
@@
Let $u\in C^{4}(\Omega)\cap C^{3}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{3}$ domain $\Omega\subset\mathbb{R}^{n}$. Suppose that $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ satisfy the boundary condition (1.2) with $G_{pp}\leq 0$ and $f^{\frac{1}{k-l}}(x,z)\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$. Then for any tangential vector field $\tau$ with $|\tau|\leq 1$ and constant $\varepsilon>0$, we have
@@
where $C_{\varepsilon}$ is a positive constant depending on $\varepsilon$, $\overline{\Omega}$, $G$, $n$, $k$, $l$, $\gamma$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $\beta_{0}$, $|u|_{C^{1}}$.
@@
(4.11) $v_{\tau}=u_{\tau\tau}+\frac{1}{2}|u_{\tau}|^{2}$
attains a maximum over $\partial\Omega$ and tangential vector $\tau$ satisfied $|\tau|\leq 1$, at $y_{0}\in\partial\Omega$ and vector $\tau=\tau_{0}$. Assume $y_{0}=0$ and $\tau_{0}=e_{1}=(1,0,\cdots,0)$. Set $b=\frac{\nu_{1}}{\left\langle\beta,\nu\right\rangle}$ and $\tau=e_{1}-b\beta$, then, at any point on $\partial\Omega$
(4.12) $v_{1}=v_{\tau}+b\left(2u_{\tau\beta}+u_{\beta}u_{\tau}\right)+b^{2}\left(u_{\beta\beta}+\frac{1}{2}u_{\beta}^{2}\right)$
satisfy $v_{1}(0)=v_{\tau}(0)$, $b(0)=0$ and $\tau(0)=e_{1}$. Let $h=\frac{1}{\left\langle\beta,\nu\right\rangle}(2u_{\tau\beta}+u_{\beta}u_{\tau})$, then
(4.13) $\left|h-h(0)\right|\leq C(1+M_{2})|x|,\quad\text{on}\;\partial\Omega,$
@@
where $C_{1}$ is a positive constant depending on $G$, $\beta_{0}$, $\overline{\Omega}$, $|u|_{C^{1}}$. Thus, we set
(4.15) $\widetilde{v}_{1}:=v_{1}-h(0)\nu_{1}-C_{1}(1+M_{2})|x|^{2},$
@@
here $Q$ is a non-negative function in $C^{2}(\overline{\Omega})$ satisfying $Q\geq|\tau|^{2}$ on $\partial\Omega$ and $Q(0)=1$. Remark that
@@
Assume $v_{1}(0)>0$ and $C_{1}$ large enough such that $Q>0$ in $\overline{\Omega}$. Now differentiating $G(x,z,p)=0$ twice in direction $\tau$ with $\tau(0)=e_{1}$ and using $G_{pp}\leq 0$, we obtain
(4.18) $D_{\beta}u_{11}(0)\geq\;-C\left(1+M_{2}\right).$
Next, we consider a new function
(4.19) $\overline{v}=\widetilde{v}_{1}-v_{1}(0)Q-K(1+M_{2})\phi,$
where $\phi\in C^{2}(\overline{\Omega})$ is a negative define function in $\Omega$ satisfying $\phi=0$ on $\partial\Omega$, $\phi_{\nu}=-1$ and $\phi_{\tau}=0$ on $\partial\Omega$, $K$ is a sufficiently large constant to be determined later. At $y_{0}=0$,
(4.20) $\begin{split}D_{\beta}\overline{v}=&\;D_{\beta}\left[\widetilde{v}_{1}-v_{1}(0)Q\right]-K(1+M_{2})\phi_{\beta}>0\end{split}$
with $K>\frac{C_{1}}{\beta_{0}}$. The specific details can be referred to Jiang-Trudinger [21, 23].
Since $\phi=0$ on $\partial\Omega$ and $\widetilde{v}_{1}\leq v_{1}(0)Q$, then $\overline{v}\leq 0\;\text{on}\;\partial\Omega$ and $\overline{v}(0)=0$. Hence $\overline{v}$ takes its maximum at $\overline{y}_{0}\in\Omega$. We suppose that the matrix $u_{ij}(\overline{y}_{0})$ is diagonal. According to (4.5) and Lemma 2.5, we have
(4.21) $0\geq L\overline{v}(\overline{y}_{0})=$ $\;F^{ij}u_{11ij}+F^{ij}u_{1i}u_{1j}+u_{1}F^{ij}u_{1ij}$
$\;-h(0)F^{ij}\nu_{1ij}-2C_{1}(1+M_{2})\sum_{i=1}^{n}F^{ii}$
$\geq$ $\;-C(1+M_{2})\sum_{i=1}^{n}F^{ii}+Cu_{11}^{2}\sum_{i=1}^{n}F^{ii},$
then $u_{11}(\overline{y}_{0})\leq\sqrt{C(1+M_{2})}\leq\varepsilon M_{2}+C_{\varepsilon}$, where $\varepsilon$ is any positive constant and $C_{\varepsilon}$ depends on $\varepsilon$, $n$, $k$, $l$, $\gamma$, $\overline{\Omega}$, $G$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $\beta_{0}$, $|u|_{C^{1}}$. Set $\phi\geq-\frac{\varepsilon}{4K}$ and $Q>\frac{1}{2}$, we get
@@
Since $v_{\tau}$ attains the maximum at 00 and $\tau=e_{1}$ on $\partial\Omega$, then $v_{\tau}\leq v_{1}(0)\leq\varepsilon M_{2}+C_{\varepsilon}$, which implies that
@@
Let $u\in C^{4}(\Omega)\cap C^{3}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{3}$ domain $\Omega\subset\mathbb{R}^{n}$. Suppose that $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ satisfy the boundary condition (1.2) with $G_{pp}=0$ and $f^{\frac{1}{k-l}}(x,z)\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$. Then for any $\varepsilon>0$,
@@
where $C_{\varepsilon}$ is a positive constant depending on $\varepsilon$, $\overline{\Omega}$, $G$, $n$, $k$, $l$, $\gamma$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $\beta_{0}$, $|u|_{C^{1}}$.
@@
(4.25) $\overline{V}=G(x,z,p)+\frac{a_{1}}{2}|Du-Du(x_{0})|^{2},$
here $a_{1}\leq 1$ is a positive constant. Since $\widetilde{F}$ is homogeneous of degree $1$ and $G_{pp}=0$, we have
@@
$\;+a_{1}\sum_{k=1}^{n}\widetilde{F}^{ij}u_{ik}u_{jk}+a_{1}\sum_{k=1}^{n}(u_{k}-u_{k}(x_{0}))\widetilde{F}^{ij}u_{ijk}$
$\geq$ $\;-C\sum_{i=1}^{n}\widetilde{F}^{ii}+a_{1}\sum_{k=1}^{n}\widetilde{F}^{ij}u_{ik}u_{jk}+2\widetilde{F}^{ij}G_{x_{i}p_{k}}u_{jk}+2\widetilde{F}^{ij}G_{zp_{k}}u_{i}u_{jk}$
@@
where $\widetilde{L}=\widetilde{F}^{ij}\partial_{ij}$ and the last inequality used $\sum_{i=1}^{n}\widetilde{F}^{ii}\geq C(n,k,l,\gamma)>0$.
@@
(4.27) $\overline{\phi}=-c\phi+\frac{a_{2}}{2}|x-x_{0}|^{2},\quad\quad\text{in}\;\Omega_{\mu},$
with
(4.28) $\phi=-d+K_{3}d^{2},$
@@
(4.29) $\widetilde{L}\overline{\phi}=-c\widetilde{F}^{ij}\phi_{ij}+a_{2}\widetilde{F}^{ij}\delta_{ij}\leq(a_{2}-c\delta)\sum_{i=1}^{n}\widetilde{F}^{ii},\quad\text{in}\;\Omega_{\mu},$
provided $\mu\leq\frac{1}{4K_{3}}$, where $\delta>0$ small enough and $K_{3}$ large enough depending on the curvature of $\partial\Omega$, $n$, $k$, $l$. Let $c\geq\frac{a_{2}}{\delta}+\frac{C}{\delta}\left(1+\frac{1}{a_{1}}\right)$, we obtain
(4.30) $\widetilde{L}\left(\overline{V}-\overline{\phi}\right)\geq 0,\quad\text{in}\;\Omega_{\mu}.$
Next, we deal with $\overline{V}$ and $\overline{\phi}$ on the boundary of $\Omega_{\mu}$.
$\mathbf{Case\;\mathrm{\mathbf{I}}}$: For any $x\in\partial\Omega$, $\overline{V}=\frac{a_{1}}{2}|Du-Du(x_{0})|^{2}$ and $\overline{\phi}=\frac{a_{2}}{2}|x-x_{0}|^{2}$, then
(4.31) $\begin{split}|Du-Du(x_{0})|\leq C(1+M_{2}^{\prime})|x-x_{0}|.\end{split}$
For any $y\in\Omega_{\mu}$ and $y^{\prime}$ the closest point on $\partial\Omega$, we obtain
(4.32) $\begin{split}|Du(y)-Du(x_{0})|^{2}\leq C(1+(M_{2}^{\prime})^{2}+M_{2})\left(|y^{\prime}-x_{0}|^{2}+d\right).\end{split}$
Choose $a_{2}=a_{1}(1+(M_{2}^{\prime})^{2})+M_{2})$ and $c\geq a_{2}$ such that
@@
$\mathbf{Case\;\mathrm{\mathbf{II}}}$: On $\partial\Omega_{\mu}\cap\Omega$, by choosing $c\geq\frac{C}{\mu}$, we have
(4.34) $\overline{V}-\overline{\phi}\leq 0,\quad\quad\quad\text{on}\;\partial\Omega_{\mu}\cap\Omega.$
Therefore,
(4.35) $\overline{V}-\overline{\phi}\leq 0,\quad\quad\quad\text{in}\;\Omega_{\mu}.$
Since $\overline{V}(x_{0})=\overline{\phi}(x_{0})=0$,
(4.36) $D_{\beta}\overline{V}(x_{0})\leq D_{\beta}\overline{\phi}(x_{0}),$
@@
(4.37) $u_{\beta\beta}(x_{0})\leq c\sup_{\partial\Omega}\left\langle\nu,\beta\right\rangle+C.$
Fix the constant $c$ such that
(4.38) $c\leq\frac{a_{2}+C(1+\frac{1}{a_{2}})}{\delta}+a_{2}+\frac{C}{\mu}\leq\frac{C}{\delta}\varepsilon_{1}M_{2}+C_{\varepsilon_{1}}(1+M_{2}^{\prime}),$
where $C$ depends on $G$, $n$, $k$, $l$, $\gamma$, $\overline{\Omega}$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $|u|_{C^{1}}$ and $\beta_{0}$. For any $\varepsilon>0$, taking $a_{1}=\frac{1}{1+\varepsilon_{1}M_{2}}$ and $\varepsilon_{1}\leq\frac{\varepsilon\delta}{C\sup_{\partial\Omega}\left\langle\nu,\beta\right\rangle}$ for a further constant $C$ in (4.38), we get
@@
Thus, Lemma 4.4 can be obtained by the arbitrariness of $x_{0}$ and Lemma 4.3. ∎
We then directly arrive at the following $C^{2}$ estimates.
Theorem 4.5.
Let $u\in C^{4}(\Omega)\cap C^{3}(\overline{\Omega})$ be an admissible solution of the boundary value problem (1.1) in a bounded $C^{3}$ domain $\Omega\subset\mathbb{R}^{n}$. Suppose that $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ satisfy the boundary condition (1.2) with $G_{pp}=0$ and $f^{\frac{1}{k-l}}(x,z)\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$. Then
@@
where $C$ is a positive constant depending on $\overline{\Omega}$, $G$, $n$, $k$, $l$, $\gamma$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $\beta_{0}$, $|u|_{C^{1}}$.
@@
Combining Lemma 4.1-Lemma 4.4, we obtain
(4.41) $\sup_{\partial\Omega}|u_{\xi\xi}|\leq\varepsilon M_{2}+C_{\varepsilon},$
for any constant unit vector $\xi$ and constant $\varepsilon>0$. Applying global second order derivative estimates, we get
(4.42) $\sup_{\overline{\Omega}}|D^{2}u|\leq C.$
Finally, we complete the proof of Theorem 1.3. From the maximal principle, $C^{0}$ estimates can be directly obtained by the existence of the admissible subsolution and $f_{z}\geq 0,\ \varphi_{z}>0$. Then based on the a priori estimates for non-degenerate case, $C^{2,\alpha}\;(0<\alpha<1)$ estimates can be established by Evans-Krylov theory and higher order estimates followed by Schauder theory. Then the existence result can be derived by the continuity method and the uniqueness assertion is immediate from the maximum principle, more details see [15]. Theorem 1.3 can be proved by approximation as in [25].

### v2

@@ 4.1. Global second order estimates
In this subsection, we establish global second order estimates for admissible solutions of problem (1.1). For the degenerate Hessian quotient type equations, the regularity condition $f^{\frac{1}{k-l}}\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$ is imposed to ensure that the estimate remain independent of $\inf_{\overline{\Omega}}f$.
@@
Let $u\in C^{4}(\Omega)$ be an admissible solution of problem (1.1) in a bounded $C^{4}$ domain $\Omega\subset\mathbb{R}^{n}$. Suppose $0\leq f^{\frac{1}{k-l}}\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$, then there exists a positive constant $C$ depending on $\overline{\Omega}$, $n$, $k$, $l$, $\gamma$, $|u|_{C^{1}}$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$ s.t.
@@
Consider an auxiliary function for a vector $\xi\in\mathbb{R}^{n}$
@@
Assume that $V$ takes its maximum at an interior point $x_{0}\in\Omega$ and a unit vector $\xi_{0}$. Without loss of generality, we can choose the coordinate system $e_{1},e_{2},\cdots,e_{n}$ at $x_{0}$ such that $e_{1}(x_{0})=\xi_{0}$ and $\{u_{ij}(x_{0})\}$ is diagonal with $\max u_{ii}(x_{0})=u_{11}(x_{0})$. Set
@@
Recall the definition $\widetilde{f}=f^{\frac{1}{k-l}}$ and use (2.5), Lemma 2.4 (4)
(4.5) $\partial_{11}f=$ $\,(k-l)f^{1-\frac{1}{k-l}}\partial_{11}\widetilde{f}++\left(1-\frac{1}{k-l}\right)\frac{(\partial_{1}f)^{2}}{f}$
$\geq$ $\,-Cf^{1-\frac{1}{k-l}}-Cu_{11}f^{1-\frac{1}{k-l}}+\left(1-\frac{1}{k-l}\right)\frac{(\partial_{1}f)^{2}}{f}$
$\geq$ $\,-C\sum_{i=1}^{n}F^{ii}-Cu_{11}\sum_{i=1}^{n}F^{ii}+\left(1-\frac{1}{k-l}\right)\frac{(\partial_{1}f)^{2}}{f}.$
By (2.4), (2.6), (4.4), (4.5) and the condition $f^{\frac{1}{k-l}}\in C^{1,1}$
(4.6) $\begin{split}0\geq F^{ij}\widetilde{V}_{ij}=&\;F^{ij}u_{11ij}+F^{ij}u_{1i}u_{1j}+F^{ij}u_{1}u_{1ij}\\ \geq&\,\partial_{11}f-F^{ij,rs}u_{ij1}u_{rs1}+Cu_{11}^{2}\sum_{i=1}^{n}F^{ii}+u_{1}\partial_{1}f\\ \geq&\;-C\sum_{i=1}^{n}F^{ii}-Cu_{11}\sum_{i=1}^{n}F^{ii}+u_{11}^{2}\sum_{i=1}^{n}F^{ii},\end{split}$
@@ 4.2. Boundary second order estimates
In this subsection, we establish boundary second order estimates, and divide the proof into three cases: double tangential derivatives, mixed tangential-oblique derivatives and double oblique derivatives. We assume that the function $G(x,u,Du)$ and the unit inner normal vector field $\nu$ have been extended to $\overline{\Omega}$, to be constant along normals to $\partial\Omega$ in some neighbourhood $\Omega_{0}$ of $\partial\Omega$.
@@
Let $u\in C^{2}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{2}$ domain $\Omega\subset\mathbb{R}^{n}$, the oblique boundary condition $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$. Then for any tangential vector field $\tau$
@@
where $\beta=G_{p}(x,u,Du)$ and the positive constant $C$ depends on $\overline{\Omega},|G|_{C^{1}},|u|_{C^{1}}$.
@@
Differentiating $G(x,u,Du)=0$ in problem (1.1) with respect to $\tau$, we have
@@
Let $u\in C^{4}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{4}$ domain $\Omega\subset\mathbb{R}^{n}$. Suppose that the oblique boundary condition $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ satisfies (1.2) with $G_{pp}\leq 0$ and $0\leq f^{\frac{1}{k-l}}\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$. Then for any tangential vector field $\tau$ with $|\tau|\leq 1$ and constant $\varepsilon>0$, we have
@@
where $C_{\varepsilon}$ is a positive constant depending on $\varepsilon$, $\overline{\Omega}$, $n$, $k$, $l$, $\gamma$, $\beta_{0}$, $|u|_{C^{1}}$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$ and $|G|_{C^{2}}$.
@@
$v_{\tau}=u_{\tau\tau}+\frac{1}{2}|u_{\tau}|^{2}$
attains a maximum at $y_{0}\in\partial\Omega,\tau=\tau_{0}$. Without loss of generality, we assume $y_{0}=0$ and $\tau_{0}=e_{1}=(1,0,\cdots,0)$. Set $b=\frac{\nu_{1}}{\left\langle\beta,\nu\right\rangle}$ and $\tau=e_{1}-b\beta$, thus at any point on $\partial\Omega$
$v_{1}=v_{\tau}+b\left(2u_{\tau\beta}+u_{\beta}u_{\tau}\right)+b^{2}\left(u_{\beta\beta}+\frac{1}{2}u_{\beta}^{2}\right),$
and $v_{1}(0)=v_{\tau}(0)$, $b(0)=0$, $\tau(0)=e_{1}$. Let $h=\frac{1}{\left\langle\beta,\nu\right\rangle}(2u_{\tau\beta}+u_{\beta}u_{\tau})$, then on $\partial\Omega$
$\left|h(x)-h(0)\right|\leq C(1+M_{2})|x|,$
@@
where $C_{1}$ is a positive constant depending on $\beta_{0}$, $\overline{\Omega}$, $|u|_{C^{1}}$, $|G|_{C^{2}}$.
Next, we consider
$\widetilde{v}_{1}=v_{1}-h(0)\nu_{1}-C_{1}(1+M_{2})|x|^{2},$
@@
$Q$ is a non-negative function in $C^{2}(\overline{\Omega})$ satisfying $Q\geq|\tau|^{2}$ on $\partial\Omega$ and $Q(0)=1$. Remark that
@@
Finally, we choose a new auxiliary function
$\overline{v}=\widetilde{v}_{1}-v_{1}(0)Q-K(1+M_{2})\phi,$
where $\phi\in C^{2}(\overline{\Omega})$ is a function in $\Omega$ satisfying $\phi=0$, $\phi_{\nu}=-1$ and $\phi_{\tau}=0$ on $\partial\Omega$, $K$ is a sufficiently large constant. Now differentiating $G(x,u,Du)=0$ twice in direction $\tau$ with $\tau(0)=e_{1}$ and using $G_{pp}\leq 0$, we obtain
$D_{\beta}u_{11}(0)\geq\;-C\left(1+M_{2}\right).$
At $y_{0}=0$
$D_{\beta}\overline{v}=D_{\beta}\left[\widetilde{v}_{1}-v_{1}(0)Q\right]-K(1+M_{2})\phi_{\beta}>0$
with $K>\frac{C_{1}}{\beta_{0}}$. The specific details can be referred to Jiang-Trudinger [24, 26]. On $\partial\Omega$ $\phi=0$, from (4.12)
$\overline{v}(0)=0,\quad\overline{v}\leq 0,\,\text{on}\,\partial\Omega,$
hence $\overline{v}$ takes its maximum at $\overline{y}_{0}\in\Omega$. Suppose that the matrix $\{u_{ij}(\overline{y}_{0})\}$ is diagonal, at $\overline{y}_{0}$
$0\geq L\overline{v}=$ $\;F^{ij}u_{11ij}+F^{ij}u_{1i}u_{1j}+u_{1}F^{ij}u_{1ij}-h(0)F^{ij}\nu_{1ij}-2C_{1}(1+M_{2})\sum_{i=1}^{n}F^{ii}$
$\;+2\frac{\beta_{1}}{\left\langle\beta,\nu\right\rangle}(0)v_{1}(0)F^{ij}\nu_{1ij}-2C_{1}v_{1}(0)\sum_{i=1}^{n}F^{ii}-N(1+M_{2})F^{ij}\phi_{ij}$
(4.13) $\geq$ $\;-C(1+M_{2})\sum_{i=1}^{n}F^{ii}+Cu_{11}^{2}\sum_{i=1}^{n}F^{ii},$
this implies $u_{11}(\overline{y}_{0})\leq\sqrt{C(1+M_{2})}\leq\varepsilon M_{2}+C_{\varepsilon}$, where $\varepsilon$ is a positive constant and $C_{\varepsilon}$ depends on $\varepsilon$, $n$, $k$, $l$, $\gamma$, $\beta_{0}$, $\overline{\Omega}$, $|u|_{C^{1}}$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $|G|_{C^{2}}$. Choosing $\phi\geq-\frac{\varepsilon}{4K}$ and $Q>\frac{1}{2}$, we get
@@
Since $v_{\tau}$ attains the maximum at 00 and $\tau=e_{1}$ on $\partial\Omega$, then $v_{\tau}\leq v_{1}(0)\leq\varepsilon M_{2}+C_{\varepsilon}$, which yields that
@@
Now it remains to establish the double oblique derivatives on the boundary.
@@
Let $u\in C^{3}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{3}$ domain $\Omega\subset\mathbb{R}^{n}$. Suppose that the oblique boundary condition $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ satisfies (1.2) with $G_{pp}=0$ and $0\leq f^{\frac{1}{k-l}}\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$. Then for a small constant $\varepsilon>0$
@@
where $C_{\varepsilon}$ is a positive constant depending on $\varepsilon$, $n$, $k$, $l$, $\gamma$, $\beta_{0}$, $\overline{\Omega}$, $|u|_{C^{1}}$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$ and $|G|_{C^{2}}$.
@@
(4.17) $\overline{V}=G(x,u,Du)+\frac{a_{1}}{2}|Du-Du(x_{0})|^{2},$
here $a_{1}\leq 1$ is a positive constant. Since $\widetilde{F}$ is homogeneous of degree $1$ and $G_{pp}=0$
@@
$\;+a_{1}\widetilde{F}^{ij}u_{ik}u_{jk}+a_{1}(u_{k}-u_{k}(x_{0}))\widetilde{F}^{ij}u_{ijk}$
$\geq$ $\;-C\sum_{i=1}^{n}\widetilde{F}^{ii}+a_{1}\widetilde{F}^{ij}u_{ik}u_{jk}+2\widetilde{F}^{ij}G_{x_{i}p_{k}}u_{jk}+2\widetilde{F}^{ij}G_{zp_{k}}u_{i}u_{jk}$
@@
where $\widetilde{L}=\widetilde{F}^{ij}\partial_{ij}$ and we use the Cauchy-Schwarz inequality in the last inequality.
@@
$\overline{\phi}=-c\phi+\frac{a_{2}}{2}|x-x_{0}|^{2},\quad\quad\text{in}\;\Omega_{\mu},$
@@
$\widetilde{L}\overline{\phi}=-c\widetilde{F}^{ij}\phi_{ij}+a_{2}\widetilde{F}^{ij}\delta_{ij}\leq(a_{2}-c\delta)\sum_{i=1}^{n}\widetilde{F}^{ii},\quad\text{in}\;\Omega_{\mu},$
then we obtain $\widetilde{L}\left(\overline{V}-\overline{\phi}\right)\geq 0$ in $v$ by choosing $c\geq\frac{a_{2}}{\delta}+\frac{C}{\delta}\left(1+\frac{1}{a_{1}}\right)$.
Next, we deal with $\overline{V}$ and $\overline{\phi}$ on the boundary of $\Omega_{\mu}$. For any $x\in\partial\Omega$
$\overline{V}=\frac{a_{1}}{2}|Du(x)-Du(x_{0})|^{2},\quad\overline{\phi}=\frac{a_{2}}{2}|x-x_{0}|^{2},$
thus
(4.19) $\begin{split}|Du(x)-Du(x_{0})|\leq C(1+M_{2}^{\prime})|x-x_{0}|,\end{split}$
consequently
(4.20) $\begin{split}|Du(y)-Du(x_{0})|^{2}\leq C(1+(M_{2}^{\prime})^{2}+M_{2})\left(|y^{\prime}-x_{0}|^{2}+d\right),\end{split}$
where $y\in\Omega_{\mu}$ and $y^{\prime}$ is the closest point on $\partial\Omega$. Choose $a_{2}=a_{1}(1+(M_{2}^{\prime})^{2})+M_{2})$ and $c\geq a_{2}$ s.t.
@@
On $\partial\Omega_{\mu}\cap\Omega$, we have $\overline{V}-\overline{\phi}\leq 0$ if $c\geq\frac{C}{\mu}$. Therefore, $\overline{V}-\overline{\phi}\leq 0$ in $\Omega_{\mu}$. Since $\overline{V}(x_{0})=\overline{\phi}(x_{0})=0$
$D_{\beta}\overline{V}(x_{0})\leq D_{\beta}\overline{\phi}(x_{0}),$
@@
$u_{\beta\beta}(x_{0})\leq c\sup_{\partial\Omega}\left\langle\nu,\beta\right\rangle+C.$
Finally, we fix the constant $c$
(4.22) $c\leq\frac{a_{2}+C(1+\frac{1}{a_{1}})}{\delta}+a_{2}+\frac{C}{\mu}\leq\frac{C}{\delta}\varepsilon_{1}M_{2}+C_{\varepsilon_{1}}(1+M_{2}^{\prime}),$
where $C$ depends on $n$, $k$, $l$, $\gamma$, $\beta_{0}$, $\overline{\Omega}$, $|u|_{C^{1}}$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$ and $|G|_{C^{2}}$. For any $\varepsilon>0$, taking $a_{1}=\frac{1}{1+\varepsilon_{1}M_{2}}$ and $\varepsilon_{1}\leq\frac{\varepsilon\delta}{C\sup_{\partial\Omega}\left\langle\nu,\beta\right\rangle}$, we get
@@
Based on the preceding lemmas, we directly arrive at the following global $C^{2}$ estimates.
Lemma 4.5.
Let $u\in C^{4}(\overline{\Omega})$ be an admissible solution of the boundary value problem (1.1) in a bounded $C^{4}$ domain $\Omega\subset\mathbb{R}^{n}$. Suppose that the oblique boundary condition $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ satisfies (1.2) with $G_{pp}=0$ and $0\leq f^{\frac{1}{k-l}}\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$. Then
@@
where $C$ is a positive constant depending on $\overline{\Omega}$, $n$, $k$, $l$, $\gamma$, $\beta_{0}$, $|u|_{C^{1}}$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $|G|_{C^{2}}$.
@@
Combine Lemma 4.1-Lemma 4.4
$\sup_{\partial\Omega}u_{\beta\beta}\leq\varepsilon_{3}M_{2}+C_{\varepsilon_{3}}(1+\varepsilon_{1}M_{2}+C_{\varepsilon_{1}})\leq\varepsilon M_{2}+C_{\varepsilon},$
and $\sup_{\partial\Omega}|D^{2}u|\leq\varepsilon M_{2}+C_{\varepsilon}$. Then
$M_{2}\leq\sup_{\partial\Omega}|D^{2}u|+C\leq\varepsilon M_{2}+C_{\varepsilon},$
which implies $\sup_{\overline{\Omega}}|D^{2}u|\leq C$. ∎
Finally, we complete the proof of Theorem 1.3. From the maximal principle, $C^{0}$ estimates can be directly obtained by the existence of the admissible subsolution and $f_{u}\geq 0,\ \varphi_{u}>0$. Then based on the a priori estimates for non-degenerate case, $C^{2,\alpha}\;(0<\alpha<1)$ estimates can be established by Evans-Krylov theory and higher order estimates followed by Schauder theory. Then the existence result can be derived by the continuity method and the uniqueness assertion is immediate from the maximum principle, more details see [18]. Theorem 1.3 can be proved by approximation as in [28].